OpenStudy (anonymous):

a natural number when increased by 12 becomes equal to 160 times the reciprocal of given natural number . find the number

OpenStudy (gorv):

let natural number=x

OpenStudy (gorv):

reciprocal=1/x

OpenStudy (gorv):

inc by 12 x+12=160*(1/x)

OpenStudy (anonymous):

but then isn't it x+12

OpenStudy (ganpat):

soo.. the equation would be say the natural number be 'x' x + 12 = 160/x...

OpenStudy (gorv):

multiply both side by x

OpenStudy (ganpat):

can u solve further ??

OpenStudy (gorv):

x(x+12)=160

OpenStudy (gorv):

Q says after increment it equals to 160 time the reciprocal of number and 1/x is reciprocal of the number

OpenStudy (gorv):

x^2+12x-160=0

OpenStudy (anonymous):

so rhs is 160/x or 160/x+12?

OpenStudy (gorv):

x^2+20x-8x-160=0 solve it now

OpenStudy (gorv):

when we add 12 to x x+12 it become equal to 160*(1/x) x+12=160/x

OpenStudy (anonymous):

ok ill try from here thanks

OpenStudy (anonymous):

is the answer x=8 and x+12=20

OpenStudy (anonymous):

??

OpenStudy (ganpat):

x2 + 12x - 160 = 0 x2 + 20x - 8x - 160 = 0 (x+20) (x-8)= 0 x = -20 and x = 8